Enter your answer in the provided box. What is the pH of the buffer 0.13 M Na2HPO4/0.19 M KH2PO4?
Answer #1

Buffer solution contain weak acid H2PO4- and its conjugate base HPO42-. pH of buffer solution is calculated by using Henderson's equation, pH = pKa + log [conjugate base ] / [ acid ]

\therefore pH = pKa + log [HPO42-] / [ H2PO4- ]

\therefore pH = 7.198 + log 0.13 / 0.19

pH = 7.198 - 0.165

pH = 7.033

ANSWER : pH of buffer solution = 7.03